Fit the parabola is Y=a X^2+bX+c for the data x: 1, 3, 4 , 6 ,8 ,9, 11, 14 and y: 1,2, 4, 4, 5, 7, 8, 9.


1
Expert's answer
2022-05-13T02:26:05-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c}
 & X & Y & XY & X^2 & X^3 & X^4 & X^2Y\\ \hline
 & 1 & 1 & 1 & 1 & 1 & 1 & 1\\
 \hdashline
 & 3 & 2 & 6 & 9 & 27 & 81 & 18 \\
 \hdashline
 & 4 & 4 & 16 & 16 & 64 & 256 & 64 \\
 \hdashline
 & 6 & 4 & 24 & 36 & 216 & 1296 & 144 \\
 \hdashline
 & 8 & 5 & 40 & 64 & 512 & 4096 & 320 \\
 \hdashline
 & 9 & 7 & 63 & 81 & 729 & 6561 & 567 \\
 \hdashline
 & 11 & 8 & 88 & 121 & 1331 & 14641 & 968 \\
 \hdashline
 & 14 & 9 & 126 & 196 & 2744 & 38416 & 1764 \\
 \hdashline
 Sum= & 56 & 40 & 364 & 524 & 5624 & 65348 & 3846 \\
 \hdashline
\end{array}"


"\bar{X}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nX_i=\dfrac{56}{8}=7"

"\bar{Y}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nY_i=\dfrac{40}{8}=5"

"\bar{X^2}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nX_i^2=\dfrac{524}{8}=65.5"

"SS_{XX}=\displaystyle\sum_{i=1}^nX_i^2-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i)^2"

"=524-\dfrac{56^2}{8}=132"

"SS_{XY}=\displaystyle\sum_{i=1}^nX_iY_i-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i)(\sum_{i=1}^nY_i)"

"=364-\dfrac{56(40)}{8}=84"

"SS_{XX^2}=\displaystyle\sum_{i=1}^nX_i^3-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i)(\displaystyle\sum_{i=1}^nX_i^2)"

"=5624-\dfrac{56(524)}{8}=1956"

"SS_{X^2X^2}=\displaystyle\sum_{i=1}^nX_i^4-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i^2)^2"

"=65348-\dfrac{(524)^2}{8}=31026"

"SS_{X^2Y}=\displaystyle\sum_{i=1}^nX_i^2Y_i-\dfrac{1}{n}(\displaystyle\sum_{i=1}^nX_i^2)(\displaystyle\sum_{i=1}^nY_i)"

"=3846-\dfrac{524(40)}{8}=1226"

"b=\dfrac{84(31026)-1226(1956)}{132(31026)-(1956)^2}=0.772286"

"a=\dfrac{1226(132)-84(1956)}{132(31026)-(1956)^2}=-0.009173"

"c=5-0.772286(7)+0.009173(65.5)=0.194830"

"\hat{y}=-0.009173x^2+0.772286x+0.194830"